# Bayesian Perspective for Orientation Determination in Cryo-EM with Application to Structural Heterogeneity Analysis

### v2

@@ 1 Introduction
Determining the precise three-dimensional (3D) orientation of biological molecules from their noisy two-dimensional (2D) projection images is a fundamental challenge in cryo-electron microscopy (cryo-EM) bai2015cryo ; lyumkis2019challenges ; bendory2020single . This process, known as orientation estimation, is crucial for various cryo-EM applications, including 3D reconstruction algorithms scheres2012relion ; punjani2017cryosparc , heterogeneity analysis sorzano2019survey ; toader2023methods ; donnat2022deep ; zhong2021cryodrgn , and beyond maeots2022structural . For example, Figure 1(a) illustrates the role of orientation estimation within the cryo-EM 3D reconstruction workflow.
In cryo-electron tomography (cryo-ET), orientation estimation presents an additional challenge in the form of subtomogram averaging. Notably, cryo-ET suffers from higher noise levels compared to single-particle cryo-EM due to the complex and heterogeneous nature of cellular samples and the challenges of capturing data from multiple angles within thicker specimens. Subtomogram averaging offers an effective approach to enhance the signal-to-noise ratio (SNR), ultimately resulting in the reconstruction of high-resolution structures. This technique often involves extracting multiple similar subtomograms containing the target protein complex or macromolecule from a large cryo-electron tomogram reconstructed from all available tilts (typically from $60^{\circ}$ to $60^{\circ}$), followed by aligning and averaging them zhang2019advances ; watson2024advances . Unlike traditional cryo-EM, this process typically aligns 3D structures directly, without the direct use of 2D projections (see Figure 1(b) for an illustration).
@@
The common approach to estimating the rotation of an observation in the models above involves scanning through a pre-defined set of possible rotations and selecting the one that either maximizes the correlation or minimizes the distance to the given reference (2D projection image or 3D molecular structure, depending on the application). Typically, these metrics involve weighted correlations and distances, where the weights account for the noise characteristics scheres2012relion .
@@
From an estimation theory perspective, this process corresponds to the maximum likelihood estimator (MLE), which does not incorporate prior information about the rotation distribution, i.e., the distribution of the 3D rotation $g$ appearing in models (1.1) or (1.2). A natural extension is the maximum a posteriori (MAP) estimator, which combines the data likelihood with a prior distribution on rotations. Notably, when a uniform prior is assumed, indicating that all orientations are equally likely a priori, the MAP estimator reduces to the MLE estimator steven1993fundamentals . A more rigorous treatment can be found in Section 2.3.
@@
However, Bayesian theory provides a much deeper and richer statistical framework that leads to improvement: replacing the MLE and MAP estimators with the Bayes estimator. The full potential of this estimator, that provides optimal accuracy according to a user-defined loss function and allows for integrating prior knowledge about rotation distribution, remained untapped so far.
@@
The term rotation distribution refers to the probabilistic law governing how orientations are distributed over all possible rotations in 3D space. In cryo-EM, this concept is critical because molecules are first suspended in a thin aqueous solution and subsequently vitrified into amorphous ice, at which point their orientations become fixed but remain uncontrolled. Ideally, all orientations would be equally likely, corresponding to a uniform distribution. In practice, however, molecules often adopt preferred orientations, typically due to interactions with the air–water interface or other sample-preparation factors tan2017addressing ; lyumkis2019challenges ; li2021effect . As a result, certain orientations are disproportionately represented, introducing systematic biases that can degrade reconstruction quality and complicate orientation estimation. Accounting for these deviations from uniformity is therefore essential for both methodological development and the interpretation of cryo-EM data.
@@
Currently, the most convenient way to obtain orientation estimates is through modules of reconstruction software such as RELION scheres2012relion and cryoSPARC punjani2017cryosparc , which output the MLE orientation estimator described earlier. These tools are primarily designed for 3D structure reconstruction and therefore treat orientation estimation as a nuisance problem; nevertheless, researchers frequently use the reported orientation estimates for downstream tasks such as heterogeneity analysis gilles2025cryo ; zhong2021cryodrgn and structural validation rosenthal2003optimal . However, relying on MLE-based estimates in these contexts can potentially lead to suboptimal results, especially under low-SNR conditions where MLE estimators are known to perform poorly. This work therefore explores opportunities to improve orientation estimation by exploiting Bayesian methods that incorporate prior knowledge of the rotation distribution.
@@
The Bayesian framework has become a powerful and widely adopted tool in cryo-EM, now recognized as the leading method for recovering 3D molecular structures scheres2012relion ; punjani2017cryosparc ; toader2023methods ; gilles2025cryo . It effectively addresses challenges like overfitting and parameter tuning while enhancing interpretability scheres2012bayesian . By explicitly modeling uncertainties, Bayesian methods enable more accurate and robust reconstructions of molecular structures, driving significant advances in both resolution and structural flexibility. While the majority of these methods focus on the task of structure reconstruction (e.g., see Section 4) and aim to achieve the MAP estimator for the volume structure scheres2012relion ; punjani2017cryosparc , the Bayesian framework offers broader possibilities for addressing other problems.
@@
One prominent example is heterogeneity analysis, where the goal is to capture structural variability that may arise from differences in composition, discrete states, or continuous conformational changes. A substantial portion of state-of-the-art methods gilles2025cryo ; zhong2021cryodrgn ; levy2024end ; luo2023opus ; punjani2017cryosparc ; punjani20213d make the simplifying assumption that particle poses are already known and fixed, typically obtained from a consensus refinement procedure; these pose estimates then form the basis for subsequent modeling of structural variability. In this work, we consider continuous conformational heterogeneity and demonstrate in Section 5 that the quality of orientation estimation has a direct impact on the fidelity of the recovered conformational landscapes.
Another important application is 3D volume alignment. Unlike the setting of (1.2), where an observation is aligned to a known reference volume, this task involves aligning two noisy volumes of the same structure with unknown relative rotations. This scenario arises, for instance, in the computation of the Fourier Shell Correlation (FSC) curve, a standard tool for estimating the spectral signal-to-noise ratio (SSNR) and resolution. As shown in Appendix B, this problem can be expressed within the same statistical framework as (1.2), differing only by an additional noise term, which highlights its close methodological connection to orientation estimation.
A further noteworthy application concerns validation. A classic example is the method of tilt pairs introduced by Rosenthal and Henderson rosenthal2003optimal , in which a small subset of images (e.g., around ten) is collected at known relative tilt angles. While the absolute particle orientations remain unknown, the relative orientations between tilt pairs are specified by the experimental geometry. The degree of agreement between estimated and known relative rotations then provides a direct validation of orientation assignment accuracy.
@@
In this work, we introduce a versatile Bayesian framework for orientation estimation with strong statistical guarantees and high flexibility. Section 2 formulates the problem and develops the Bayesian MMSE estimator. We show theoretically that the MMSE estimator coincides with the MLE estimator in high-SNR regimes (Proposition 2.2), while consistently outperforming it in low-SNR conditions (Figure 1), a typical scenario in cryo-EM and cryo-ET bendory2020single .
@@
In Section 4, we connect the MMSE estimator to 2D image recovery and 3D structure reconstruction, showing that it naturally leads to the expectation-maximization (EM) framework for models without projections (Proposition 4.1). Empirical results confirm that the MMSE estimator consistently outperforms the MLE and MAP estimators in both tasks, providing higher accuracy and greater robustness against the “Einstein from Noise” artifact henderson2013avoiding ; balanov2024einstein (Figures 45).
@@
Beyond reconstruction accuracy, we demonstrate that orientation estimation quality has a direct and substantial impact on downstream structural heterogeneity analysis. Understanding structural variability is essential for characterizing the dynamic behavior of macromolecular complexes. Methods such as RECOVAR gilles2025cryo , designed to recover structural heterogeneity, typically assume known particle poses (i.e., fixed-pose methods). In reality, these poses must be inferred, and MLE-derived estimates are commonly used in practice. We show that replacing MLE-derived poses with MMSE estimates in RECOVAR substantially improves the recovery of the latent conformational manifold, bringing results closer to those obtained using ground-truth poses. These findings establish MMSE orientation estimation not only as a tool that enhances reconstruction fidelity, but also as a facilitator that advances state-of-the-art continuous heterogeneity analysis.
@@
Figure 1: Comparison between maximum cross-correlation ($\hat{g}_{\text{MLE}}$) and Bayesian ($\hat{g}_{\text{MMSE}}$) orientation estimators in cryo-electron microscopy (cryo-EM) and cryo-electron tomography (cryo-ET) applications. The figure illustrates the general workflow in cryo-EM and cryo-ET techniques, highlighting the role of orientation estimation in each technique. Panel (a) illustrates the model with 2D projections (single-particle cryo-EM model, (1.1)), while panel (b) shows the model of the subtomogram averaging in cryo-ET, (1.2)). (a) Cryo-EM involves imaging macromolecules embedded in a thin layer of vitreous ice using an electron beam in a transmission electron microscope (TEM). The process generates 2D projection images (micrographs) of particles in unknown 3D orientations. These 2D particles are then identified and extracted from the micrographs, forming the basis for subsequent steps of the macromolecule’s 3D structure reconstruction. (b) Cryo-ET involves imaging a sample from multiple known tilt angles (typically from $60^{\circ}$ to $60^{\circ}$) to create 2D projections, which are then combined computationally to reconstruct 3D subtomograms. In this context, a subtomogram refers to a small volume containing an individual 3D particle. The subtomograms are extracted by a particle picker algorithm for further analysis. (a+b): The rotation estimation problem involves determining the relative orientation of a noisy 2D particle (in cryo-EM) or a noisy 3D subtomogram (in cryo-ET) relative to a reference volume $V$. The reference volume structure used in both setups is identical and corresponds to the 80S ribosome wong2014cryo . Under high SNR conditions, both rotation estimators closely approximate the true relative rotation. However, as the SNR decreases, the estimation accuracy deteriorates. Importantly, across all SNR levels, the geodesic angular distance between the MMSE orientation estimator and the true rotation consistently remains lower than that of the MLE orientation estimator. For (a), the estimation was conducted using a grid size of $L=3000$ samples of the rotation group $\mathsf{SO}(3)$, while for (b), a grid size of $L=300$ was used. Each point in the two curve plots represents the average error computed over 3000 trials.
@@ 2 Problem formulation and the MMSE orientation estimator
In this section, we present a particular Bayes estimator, the minimum mean square error (MMSE) estimator, for orientation determination within the Bayesian framework. We begin by introducing a flexible mathematical model that encompasses various typical applications involving orientation estimation. Following this, we present a couple of metrics designed to assess the quality of our estimators, providing a robust framework for evaluating performance. Finally, we introduce the class of Bayes estimators, with the MMSE estimator serving as a primary example.
@@ 2.1 Mathematical model for orientation estimation
We begin with a continuous-domain formulation. Let $V(x):\mathbb{R}^{n}\to\mathbb{R}$ denote a reference structure, where $n=2$ for 2D images and $n=3$ for 3D volumes. For example, when $n=3$, $V$ corresponds to the continuous 3D electron density, as commonly used in cryo-EM or cryo-ET. Let $\mathsf{G}\subset\mathrm{SO}(n)$ be a compact group of rotations correspondingly, and let $g\in\mathsf{G}$ be an unknown transformation drawn from a distribution $\Lambda$ over $\mathsf{G}$.
@@
    2D template matching. In this case, $\Pi$ is solely the discretization sampling operator, $d=N\times N$ with $N$ the grid size of the 2D image, $g\in\mathsf{G}=\mathrm{SO}(2)$ is a 2D in-plane rotation, and V{V} is a given 2D template image.
@@
    3D structure alignment in cryo-ET. In this scenario, we consider a special case of (2.1) where $\Pi$ is solely the discretization sampling operator, $d$ corresponds to the total dimension of 3D subtomograms, $g\in\mathsf{G}=\mathsf{SO}(3)$ is a 3D rotation, and V{V} is a given 3D volume. Notably, the 3D alignment problem is also a critical step in the computational pipeline of cryo-EM, see, e.g., singer2024alignment ; harpaz2023three .
@@ 2.2 Preliminaries
It is important to note that the framework is flexible and can accommodate other loss functions over any group $\mathsf{G}$. For a more comprehensive discussion of rotation metrics, we direct readers to the work of hartley2013rotation ; huynh2009metrics . Similar metrics and estimation procedures apply to the $\mathrm{SO}(2)$ estimation problem, we omit a detailed discussion for brevity.

### v3

@@ 1 Introduction
Determining the precise three-dimensional (3D) orientation of biological molecules from their noisy two-dimensional (2D) projection images is a fundamental challenge in cryo-electron microscopy (cryo-EM) Bai et al. (2015); Lyumkis (2019); Bendory et al. (2020). This process, known as orientation estimation, is crucial for various cryo-EM applications, including 3D reconstruction algorithms Scheres (2012b); Punjani et al. (2017), heterogeneity analysis Sorzano et al. (2019); Toader et al. (2023); Donnat et al. (2022); Zhong et al. (2021), and beyond Mäeots and Enchev (2022). For example, Figure 1(a) illustrates the role of orientation estimation within the cryo-EM 3D reconstruction workflow.
In cryo-electron tomography (cryo-ET), orientation estimation presents an additional challenge in the form of subtomogram averaging. Notably, cryo-ET suffers from higher noise levels compared to single-particle cryo-EM due to the complex and heterogeneous nature of cellular samples and the challenges of capturing data from multiple angles within thicker specimens. Subtomogram averaging offers an effective approach to enhance the signal-to-noise ratio (SNR), ultimately resulting in the reconstruction of high-resolution structures. This technique often involves extracting multiple similar subtomograms containing the target protein complex or macromolecule from a large cryo-electron tomogram reconstructed from all available tilts (typically from $60^{\circ}$ to $60^{\circ}$), followed by aligning and averaging them Zhang (2019); Watson and Bartesaghi (2024). Unlike traditional cryo-EM, this process typically aligns 3D structures directly, without the direct use of 2D projections (see Figure 1(b) for an illustration).
@@
The common approach to estimating the rotation of an observation in the models above involves scanning through a pre-defined set of possible rotations and selecting the one that either maximizes the correlation or minimizes the distance to the given reference (2D projection image or 3D molecular structure, depending on the application). Typically, these metrics involve weighted correlations and distances, where the weights account for the noise characteristics Scheres (2012b).
@@
From an estimation theory perspective, this process corresponds to the maximum likelihood estimator (MLE), which does not incorporate prior information about the rotation distribution, i.e., the distribution of the 3D rotation $g$ appearing in models (1.1) or (1.2). A natural extension is the maximum a posteriori (MAP) estimator, which combines the data likelihood with a prior distribution on rotations. Notably, when a uniform prior is assumed, indicating that all orientations are equally likely a priori, the MAP estimator reduces to the MLE estimator Steven (1993). A more rigorous treatment can be found in Section 2.3.
@@
However, Bayesian theory provides a much deeper and richer statistical framework that leads to improvement: replacing the MLE or MAP estimator with the Bayes estimator. The full potential of this estimator, that provides optimal accuracy according to a user-defined loss function and allows for integrating prior knowledge about rotation distribution, remained untapped so far.
@@
The term rotation distribution, used interchangeably with orientation distribution in this paper, refers to the probabilistic law governing how orientations are distributed over all possible rotations in 3D space. In cryo-EM, this concept is critical because molecules are first suspended in a thin aqueous solution and subsequently vitrified into amorphous ice, at which point their orientations become fixed but remain uncontrolled. Ideally, all orientations would be equally likely, corresponding to a uniform distribution. In practice, however, molecules often adopt preferred orientations, typically due to interactions with the air–water interface or other sample-preparation factors Tan et al. (2017); Lyumkis (2019); Li et al. (2021). As a result, certain orientations are disproportionately represented, introducing systematic biases that can degrade reconstruction quality and complicate orientation estimation. Accounting for these deviations from uniformity is therefore essential for both methodological development and the interpretation of cryo-EM data.
Currently, the most practical way to obtain orientation estimates is through modules of reconstruction software such as RELION Scheres (2012b) and cryoSPARC Punjani et al. (2017). These packages are primarily designed for 3D structure reconstruction and therefore treat pose estimation as a latent-variable subproblem. In particular, during refinement they compute posterior weights over a discrete set of candidate poses for each particle and use these weights for weighted averaging in the reconstruction step (a soft-assignment over poses). At the same time, for downstream usage and reporting, the metadata typically provides a single orientation per particle, which is commonly the MLE estimator, that is, the maximizer of the pose posterior. Thus, while these pipelines do incorporate latent-variable pose weighting internally for reconstruction, they do not typically output a Bayesian rotation estimate for each particle.
@@
In practice, researchers frequently use the reported per-particle orientation point estimates for downstream tasks such as heterogeneity analysis Gilles and Singer (2025); Zhong et al. (2021) and structural validation Rosenthal and Henderson (2003). However, relying solely on MLE-based estimates can be suboptimal, especially in low-SNR regimes where MLE estimators are known to perform poorly. This work therefore explores opportunities to improve per-particle orientation estimation by using Bayes estimators that explicitly average over pose uncertainty that incorporate prior knowledge of the rotation distribution.
@@
The Bayesian framework has become a powerful and widely adopted tool in cryo-EM, now recognized as the leading method for recovering 3D molecular structures Scheres (2012b); Punjani et al. (2017); Toader et al. (2023); Gilles and Singer (2025). It effectively addresses challenges like overfitting and parameter tuning while enhancing interpretability Scheres (2012a). By explicitly modeling uncertainties, Bayesian methods enable more accurate and robust reconstructions of molecular structures, driving significant advances in both resolution and structural flexibility. While the majority of these methods focus on the task of structure reconstruction (e.g., see Section 4) and aim to achieve the MAP estimator for the volume structure Scheres (2012b); Punjani et al. (2017), the Bayesian framework offers broader possibilities for addressing other problems.
@@
One prominent example is heterogeneity analysis, where the goal is to capture structural variability that may arise from differences in composition, discrete states, or continuous conformational changes. A substantial portion of state-of-the-art methods Gilles and Singer (2025); Zhong et al. (2021); Levy et al. (2024); Luo et al. (2023); Punjani et al. (2017); Punjani and Fleet (2021) make the simplifying assumption that particle poses are already known and fixed, typically obtained from a consensus refinement procedure; these pose estimates then form the basis for subsequent modeling of structural variability. In this work, we consider continuous conformational heterogeneity and demonstrate in Section 5 that the quality of orientation estimation has a direct impact on the fidelity of the recovered conformational landscapes.
Another important application is 3D volume alignment. Unlike the setting of (1.2), where an observation is aligned to a known reference volume, this task involves aligning two noisy volumes of the same structure with unknown relative rotations. This scenario arises, for instance, in the computation of the Fourier Shell Correlation (FSC) curve, a standard tool for estimating the spectral signal-to-noise ratio (SSNR) and resolution. As shown in Appendix C, this problem can be expressed within the same statistical framework as (1.2), differing only by an additional noise term, which highlights its close methodological connection to orientation estimation.
A further noteworthy application concerns validation. A classic example is the method of tilt pairs introduced by Rosenthal and Henderson Rosenthal and Henderson (2003), in which a small subset of images (e.g., around ten) is collected at known relative tilt angles. While the absolute particle orientations remain unknown, the relative orientations between tilt pairs are specified by the experimental geometry. The degree of agreement between estimated and known relative rotations then provides a direct validation of orientation assignment accuracy.
@@
In this work, we introduce a versatile Bayesian framework for orientation estimation with strong statistical guarantees and high flexibility. Section 2 formulates the problem and develops the Bayesian MMSE estimator. We show theoretically that the MMSE estimator coincides with the MLE estimator in high-SNR regimes (Proposition 2.2), while consistently outperforming it in low-SNR conditions (Figure 1), a typical scenario in cryo-EM and cryo-ET Bendory et al. (2020).
@@
In Section 4, we connect the MMSE estimator to 2D image recovery and 3D structure reconstruction, showing that it naturally leads to the expectation-maximization (EM) framework for models without projections (Proposition 4.1). Empirical results confirm that the MMSE estimator consistently outperforms the MLE and MAP estimators in both tasks, providing higher accuracy and greater robustness against the “Einstein from Noise” artifact Henderson (2013); Balanov et al. (2024) (Figures 45).
@@
Beyond reconstruction accuracy, we demonstrate that orientation estimation quality has a direct and substantial impact on downstream structural heterogeneity analysis. Understanding structural variability is essential for characterizing the dynamic behavior of macromolecular complexes. Methods such as RECOVAR Gilles and Singer (2025), designed to recover structural heterogeneity, typically assume known particle poses (i.e., fixed-pose methods). In reality, these poses must be inferred, and MLE-derived estimates are commonly used in practice. We show that replacing MLE-derived poses with MMSE estimates in RECOVAR substantially improves the recovery of the latent conformational manifold, bringing results closer to those obtained using ground-truth poses. These findings establish MMSE orientation estimation not only as a tool that enhances reconstruction fidelity, but also as a facilitator that advances state-of-the-art continuous heterogeneity analysis.
@@
Figure 1: Comparison between maximum cross-correlation ($\hat{g}_{\text{MLE}}$) and Bayesian ($\hat{g}_{\text{MMSE}}$) orientation estimators in cryo-electron microscopy (cryo-EM) and cryo-electron tomography (cryo-ET) applications. The figure illustrates the general workflow in cryo-EM and cryo-ET techniques, highlighting the role of orientation estimation in each technique. Panel (a) illustrates the model with 2D projections (single-particle cryo-EM model, (1.1)), while panel (b) shows the model of the subtomogram averaging in cryo-ET, (1.2)). (a) Cryo-EM involves imaging macromolecules embedded in a thin layer of vitreous ice using an electron beam in a transmission electron microscope (TEM). The process generates 2D projection images (micrographs) of particles in unknown 3D orientations. These 2D particles are then identified and extracted from the micrographs, forming the basis for subsequent steps of the macromolecule’s 3D structure reconstruction. (b) Cryo-ET involves imaging a sample from multiple known tilt angles (typically from $60^{\circ}$ to $60^{\circ}$) to create 2D projections, which are then combined computationally to reconstruct 3D subtomograms. In this context, a subtomogram refers to a small volume containing an individual 3D particle. The subtomograms are extracted by a particle picker algorithm for further analysis. (a+b): The rotation estimation problem involves determining the relative orientation of a noisy 2D particle (in cryo-EM) or a noisy 3D subtomogram (in cryo-ET) relative to a reference volume $V$. The reference volume structure used in both setups is identical and corresponds to the 80S ribosome Wong et al. (2014). Under high SNR conditions, both rotation estimators closely approximate the true relative rotation. However, as the SNR decreases, the estimation accuracy deteriorates. Importantly, across all SNR levels, the geodesic angular distance between the MMSE orientation estimator and the true rotation consistently remains lower than that of the MLE orientation estimator. For (a), the estimation was conducted using a grid size of $L=3000$ samples of the rotation group $\mathsf{SO}(3)$, while for (b), a grid size of $L=300$ was used. Each point in the two curve plots represents the average error computed over 3000 trials.
@@ 2 Problem formulation and the MMSE orientation estimator
In this section, we present a particular Bayes estimator for orientation determination under a squared-error loss, commonly known as the minimum mean-square error (MMSE) estimator. The name reflects that this estimator minimizes the expected posterior mean-squared error, and it is given by the posterior (conditional) mean rather than a pointwise maximization Kay (1993).
We begin by introducing a flexible mathematical model that encompasses various typical applications involving orientation estimation. Following this, we present a couple of metrics designed to assess the quality of our estimators, providing a robust framework for evaluating performance. Finally, we introduce the class of Bayes estimators, with the MMSE estimator serving as a primary example.
@@ 2.1 Mathematical model for orientation estimation
We begin with a continuous-domain formulation. Let $V(x):\mathbb{R}^{n}\to\mathbb{R}$ denote a reference structure, where $n=2$ for 2D images and $n=3$ for 3D volumes. For example, when $n=3$, $V$ corresponds to the continuous 3D electron density, as commonly used in cryo-EM or cryo-ET. Let $\mathsf{G}\subset\mathsf{SO}(n)$ be a compact group of rotations correspondingly, and let $g\in\mathsf{G}$ be an unknown transformation drawn from a distribution $\Lambda$ over $\mathsf{G}$.
@@
    2D template matching. In this case, $\Pi$ is solely the discretization sampling operator, $d=N\times N$ with $N$ the grid size of the 2D image, $g\in\mathsf{G}=\mathsf{SO}(2)$ is a 2D in-plane rotation, and V{V} is a given 2D template image.
@@
    3D structure alignment in cryo-ET. In this scenario, we consider a special case of (2.1) where $\Pi$ is solely the discretization sampling operator, $d$ corresponds to the total dimension of 3D subtomograms, $g\in\mathsf{G}=\mathsf{SO}(3)$ is a 3D rotation, and V{V} is a given 3D volume. Notably, the 3D alignment problem is also a critical step in the computational pipeline of cryo-EM, see, e.g., Singer and Yang (2024); Harpaz and Shkolnisky (2023).
@@ 2.2 Preliminaries
We note that the chordal and geodesic distances are closely related. In particular, if $\theta\triangleq d_{G}(g_{1},g_{2})$ denotes the geodesic rotation angle between $g_{1}$ and $g_{2}$ (2.8), then
$d_{F}(g_{1},g_{2})=\|g_{1}-g_{2}\|_{F}=2\sqrt{2}\,\sin(\theta/2)=2\sqrt{2}\,\sin\!\big(d_{G}(g_{1},g_{2})/2\big).$ (2.9)
Consequently, $d_{F}$ is a strictly increasing function of $d_{G}$ and, for small angular errors, the two are locally equivalent in the sense that $d_{F}(g_{1},g_{2})\approx\sqrt{2}\,d_{G}(g_{1},g_{2})$ (see, e.g., Hartley et al. (2013)). More generally, our framework is flexible and can accommodate other loss functions over any group $\mathsf{G}$. For a comprehensive discussion of metrics on $\mathsf{SO}(3)$ and their relationships, we refer the reader to Hartley et al. (2013); Huynh (2009). Similar considerations apply to $\mathsf{SO}(2)$; we omit a detailed discussion for brevity.